OpenStudy (anonymous):

Identify the 42nd term of an arithmetic sequence where a1 = -12 and a27 = 66

OpenStudy (mathmath333):

find \( d\) from the information

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{a_{1},a_{2},a_{3},a_{4}\cdot \cdot \cdot \cdot \infty \hspace{.33em}\\~\\ \implies a_{1},a_{1}+d,a_{1}+2d,a_{1}+3d \cdot \cdot \cdot \cdot \infty \hspace{.33em}\\~\\ a_{1}=-12 \hspace{.33em}\\~\\ a_{27}=66 \hspace{.33em}\\~\\ a_{27}=a_1+26d \hspace{.33em}\\~\\ a_{42}=a_1+41d \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

nth term of an AP is given by... \[a_{n}=a+(n-1)d\]\[a=-12\] for n=27, \[a{_{27}}=a+(27-1)d\] put the value of a and a27 \[66=-12+26d\] calculate d from this equation then for n=42, \[a_{42}=a+(42-1)d\]\[a_{42}=a+41d\] put the values of a and d and calculate a42